OpenStudy (anonymous):

Given f(x)=-3+2 and g(x) = 2x + 9 Find (go f)(x)

OpenStudy (anonymous):

(g o f)(x) = g(f(x)) = 2(-3+2) + 9 = 2(-1) + 9 = -2+9 = 7 so (g o f)(x) = 7

OpenStudy (anonymous):

f(x)=-3x+2?

OpenStudy (precal):

substitute f(x) into g(x)